OpenStudy (anonymous):

how are logarithms programmed as a function in a calculator or other programmable devices

OpenStudy (anonymous):

Generally it's done using a series

OpenStudy (anonymous):

The taylor series for example can get you the natural log of x (albeit a bit slowly): \[ln(x) = \sum_{n=1}^\infty \frac{1}{n}(x-1)^n\]

OpenStudy (anonymous):

oh wow interesting concept

OpenStudy (dumbcow):

thats what i was thinking there are more efficient programs they use which will give decent approximations

OpenStudy (anonymous):

Whoops, forgot the (-1)^{n+1}

OpenStudy (anonymous):

umm but what about square root functions do they just iterate throughtit

OpenStudy (anonymous):

until they find a match?

OpenStudy (anonymous):

\[ln(x) = \sum_{n=1}^\infty \frac{1}{n}(-1)^{n+1}(x-1)^n\]

OpenStudy (anonymous):

No, infinite series can get you good approximations for any function you like.

OpenStudy (dumbcow):

actually newtons method for finding zeroes can be used to approximate square roots

OpenStudy (anonymous):

thankyou polpak and dumbcow

OpenStudy (dumbcow):

your welcome, actually i take that back i dunno if they would use newtons method in a program because it would have to be recursive

OpenStudy (dumbcow):

prob infinite series like with logs